The sum of the ages of Roberto and Guadalupe is 88 years. 6 years ago, Roberto's age was 3 times Guadalupe's age. How old is Roberto now?
Answer by VFBundy(438)

Roberto's age now = x
Guadalupe's age now = 88 - x

"6 years ago, Roberto's age was 3 times Guadalupe's age":

x - 6 = 3(88 - x - 6)

x - 6 = 3(82 - x)

x - 6 = 246 - 3x

4x = 252

x = 63

Roberto's age now = x = 63

Double-check:
Roberto's age now = x = 63
Guadalupe's age now = 88 - x = 25

"The sum of the ages of Roberto and Guadalupe is 88 years":

63 + 25 = 88

"6 years ago, Roberto's age was 3 times Guadalupe's age":

63 - 6 = 3(25 - 6)

57 = 3(19)